OpenStudy (anonymous):

A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 32 feet and the child ride has a radius of 24 feet, what is the ratio between the volumes of each ride? 4:3 16:9 64:27 12:9

OpenStudy (amistre64):

volume of a cone is what? 1/3 pi r^2 * h right?

OpenStudy (amistre64):

with all things being equal but radius; that least us with R^2/r^2

OpenStudy (amistre64):

but similar would suggest that height aint equal; but ratioed as well. and since its linear we can prolly inerpolate it as well

OpenStudy (amistre64):

32/24 = H/h 4/3 = H/h H = 4h/3 1/3 pi R^2 H 32^2 4h/3 ------------ = ---------- 1/3 pi r^2 h 24^2 h 32 32 4 h --------- 24 24 3 h 8 8 4 ----- = 64(4)/27 3 3 3 maybe

OpenStudy (amistre64):

444/333 = 64/27 :)

OpenStudy (amistre64):

i reduced my 32s badly lol

OpenStudy (anonymous):

Lol it's okay, thank you very much :)

OpenStudy (amistre64):

yw :P